Bovine serum albumin (BSA) is a biochemically useful protein. A 0.257 gram sample of bovine serum albumin is dissolved in water to make 0.146 L of solution, and the osmotic pressure of the solution at 25 °C is found to be 0.658 mbar. Calculate the molecular mass of bovine serum albumin.
Answer #1

the osmotic pressure, (pi), using the following equation:

Where:

= 0.658 × 10-3 = 0.000658 bar

M is the molar concentration of dissolved species (units of mol/L), M = Mass / Molar mass / Vol = 0.257/ M/ 0.146

R is the ideal gas constant, R = 0.08314 L bar mol-1 K-1,
T is the temperature on the Kelvin scale, T = 298K

Molar mass of bovine serum albumin is 6628 g/mol